# MsgABD: EveryShipout initializing macros

### v1

@@ D.3 Test Accuracy versus DropNode Probability
In Equation 3.4, we noted that the expectation of sensitivity remains unchanged when using DropNode. However, these results were only in expectation. In practice, a high DropNode probability will result in poor communication between distant nodes, preventing the model from learning to effectively model LRIs. This is supported by the results in Table 2, where we observed a negative correlation between the test accuracy and DropNode probability. Moreover, DropNode was the only algorithm which recorded negative correlations on homophilic datasets. In Figure 6, we visualize these relationships, noting the stark contrast with Figure 2, particularly in the trends with homophilic datasets.
Appendix E Experiment Details
In this section, we expand on the details of the experiment in Section 4.
Descriptions of the Datasets. Cora McCallum et al. (2000) and CiteSeer Giles et al. (1998) are citation networks – their nodes represent scientific publications and an edge between two nodes indicates that one of them has cited the other. The features of each publication are represented by a binary vector, where each index indicates whether a specific word from a dictionary is present or absent. Several studies have showed that these datasets have high homophily in node labels Lim et al. (2021); Zhu et al. (2020) and that they are modelled much better by shallower networks than by deeper ones Zhou et al. (2020). Chameleon and Squirrel Rozemberczki et al. (2021) are networks of English Wikipedia web pages on the respective topics, and the edges between web pages indicate links between them. The task is to predict the average-monthly traffic on each of the web pages. Finally, TwitchDE Rozemberczki et al. (2021) is a network of Twitch users in Germany, with the edges between them representing their mutual follower relationships. The node features are embeddings of the games played by the users. The task is to predict whether the users use explicit language.
Training Hyperparameters. We standardise most of the hyperparameters across all experiments in order to isolate the effect of dropping probability. Specifically, we fix the size of the hidden representations in each layer at 64, and a linear readout layer is used to compute the node-level logits. The models are trained using the Adam optimizer Kingma & Ba (2015), with a learning rate of $3\times 10^{-3}$ and a weight decay of $5\times 10^{-4}$, for a total of 300 epochs. For GCN, we use symmetric normalization of the adjacency matrix to compute the edge weights Kipf & Welling (2017).
GAT Runs. It is quite problematic to train deep GAT models due to vanishing gradients (Dasoulas et al., 2021). Accordingly, we discard the runs where the model fails to learn, and performs just as well as a random classifier. Specifically, we compute the class distribution in each of the networks, and discard the runs where the test performance does not exceed the maximum proportion. This comes out to be $0.3021$, $0.2107$, $0.2288$, $0.2003$ and $0.6045$ for Cora, CiteSeer, Chameleon, Squirrel and TwitchDE, respectively.
E.1 Extension to GraphSAGE and ResGCN
Our exploration in this work has been mainly limited to stochastic regularization methods, including edge-dropping methods, and related ones like Dropout Srivastava et al. (2014), DropMessage Fang et al. (2023) and SkipNode Lu et al. (2024). We may follow a similar intuition to study other methods as well. For example, GraphSAGE Hamilton et al. (2017) operates by sampling a subset of neighbors for message aggregation in each step. From an over-squashing perspective, it is very similar to DropEdge and therefore, we can expect a similar decline in sensitivity between distant nodes when using GraphSAGE.
Figure 7: Influence distributions of a regular GCN and a ResGCN. Residual connections reduce the influence of nodes farther away by amplifying the influence of nodes close by.
On the other hand, residual connection reduce the influence of distant nodes by increasing the influence of nearby neighbors. For example, in the case of ResGCN, the operation $\bm{Z}^{\left(\ell+1\right)}=\hat{\bm{A}}\bm{Z}^{\left(\ell\right)}\bm{W}^{\left(\ell+1\right)}+\bm{Z}^{\left(\ell\right)}$ first collects information from nodes $(\ell+1)$-hops away, like in a regular GCN model, and then adds information from nodes up to $\ell$-hops away, effectively decreasing the influence of the nodes $(\ell+1)$-hops away. In other words, we may expect residual connections to be similarly unsuited for long-range tasks, not by reducing the strength of the signal over long distances, but rather by amplifying the strength of short-range signals. In Figure 7, we present the empirical influence distribution Xu et al. (2018) of a 6-layer ResGCN, computed using the Cora network. Here, we can clearly see that residual connections significantly reduce the influence of distant nodes on final representations, and can thus be expected to be unsuitable for long-range tasks.

### v2

@@ F.2 Remark on DropNode
In Equation B.15, we noted that DropNode does not suffer from loss in sensitivity. However, those results were in expectation. Moreover, our analysis did not account for the effects on the learning trajectory. In practice, a high DropNode probability would make it hard for information in the node features to reach distant nodes. This would prevent the model from learning to effectively combine information from large neighborhoods, harming generalization. In Figure 8, we visualize the relationship between test-time performance and DropNode probability. The performance monotonically decreases with increasing dropping probability, as was observed with DropEdge.
F.3 Over-squashing or Under-fitting?
Figure 9: DropEdge probability versus training accuracy of GCNs. The training performance improves with $q$, suggesting that the models are not underfitting. Instead, the reason for poor test-time performance (Figure 7) is that models are over-fitting to short-range signals during training, resulting in poor generalization.
The results in the previous subsection suggest that using random edge-dropping to regularize model training leads to poor test-time performance. We hypothesize that this occurs because the models struggle to propagate information over long distances, causing node representations to overfit to local neighborhoods. However, a confounding effect is at play: DropEdge variants reduce the generalization gap by preventing overfitting to the training set, i.e. poorer training performance. If this regularization is too strong, it could lead to underfitting, which could also explain the poor test-time performance on heterophilic datasets. This concern is particularly relevant because the heterophilic networks are much larger than homophilic ones (see Table 4), making them more prone to underfitting. To investigate this, we plot the training accuracies of deep DropEdge-GCNs on the heterophilic datasets; Figure 9 shows the results. It is clear that the models do not underfit as the dropping probability increases. In fact, somewhat unexpectedly, the training metrics improve. Together with the results in Figure 7, we conclude that DropEdge-like methods are detrimental in long-range tasks since they cause overfitting to short-range artifacts in the training data, resulting in poor generalization at test-time.
Appendix G Supplementary Experimental Results
G.1 Performance of GAT with Dropping Methods
In Table 6, we present the results of experiments in Section 5.2, but with the GAT architecture. For node classification tasks, we see the same dichotomy as in 1(a), with dropping methods significantly improving performance on homophilic networks, while being detrimental to performance on heterophilic networks. On graph classification tasks, the dropping methods improve performance, but the improvement (if any) is not statistically significant (in $28/36\approx 78\%$ cases). Note that the GAT architecture was unable to learn the Collab dataset, i.e. the performance in all cases was as good as a random classifier’s.
Table 6: Difference in mean test accuracy (%) between the best performing configuration of each dropout method and the baseline NoDrop model, with GAT as the base model. Cell colors represent p-values from a t-test evaluating whether dropout improves performance.
(a) Node classification tasks.
Dropout Homophilic Networks Heterophilic Networks
Cora CiteSeer PubMed Chameleon Squirrel TwitchDE
DropEdge $+0.483$ $+0.828$ $-0.064$ $-1.988$ $-1.125$ $-0.206$
DropNode $+0.200$ $-0.002$ $-0.181$ $-6.090$ $-2.016$ $-2.383$
DropAgg $+0.322$ $+0.797$ $+0.032$ $-7.779$ $-4.904$ $-1.014$
DropGNN $+0.519$ $+1.058$ $-0.149$ $-10.572$ $-5.214$ $-1.636$
Dropout $+0.600$ $+0.104$ $+0.265$ $-7.891$ $-2.773$ $-1.598$
DropMessage $+0.389$ $+0.039$ $+0.633$ $-1.569$ $-0.092$ $-0.139$
(b) Graph classification tasks.
Dropout Molecular Networks Social Networks
Mutag Proteins Enzymes Reddit IMDb Collab
[…]
G.2 Effect Size in Statistical Tests
Table 7: Hedges’ $g$ statistic for different dataset--model--dropout combinations. Color-coding for effect size according to Cohen [14]; red denotes negative effect, and green denotes positive effect. Medium to large positive effect sizes in bold.
X No effect    X X Small effect    X X Medium effect    X X Large effect
Dataset GNN DropEdge DropNode DropAgg DropGNN Dropout DropMessage
Cora GCN $\mathbf{+0.973}$ $\mathbf{+0.759}$ $\mathbf{+0.431}$ $\mathbf{+0.900}$ $\mathbf{+0.990}$ $+0.060$
GIN $-0.174$ $+0.205$ $+0.068$ $-0.212$ $\mathbf{+1.010}$ $\mathbf{+2.180}$
GAT $\mathbf{+0.933}$ $+0.317$ $\mathbf{+0.455}$ $\mathbf{+1.083}$ $\mathbf{+0.977}$ $\mathbf{+0.710}$
CiteSeer GCN $\mathbf{+0.848}$ $+0.268$ $\mathbf{+0.628}$ $\mathbf{+0.929}$ $+0.155$ $-0.397$
GIN $-0.254$ $\mathbf{+0.515}$ $-0.124$ $-0.854$ $+0.006$ $+0.132$
GAT $\mathbf{+0.799}$ $-0.002$ $\mathbf{+0.777}$ $\mathbf{+0.985}$ $+0.103$ $+0.040$
PubMed GCN $\mathbf{+1.039}$ $\mathbf{+1.285}$ $-0.627$ $\mathbf{+1.455}$ $\mathbf{+2.625}$ $\mathbf{+2.907}$
GIN $-0.043$ $+0.343$ $-0.546$ $-1.213$ $\mathbf{+0.423}$ $\mathbf{+1.118}$
GAT $-0.108$ $-0.320$ $+0.053$ $-0.271$ $\mathbf{+0.407}$ $\mathbf{+0.964}$
Chameleon GCN $-0.167$ $-0.174$ $-3.279$ $-0.479$ $-0.385$ $\mathbf{+0.356}$
GIN $-0.219$ $-0.567$ $-0.329$ $-0.510$ $-0.857$ $+0.290$
GAT $-0.429$ $-1.326$ $-1.945$ $-2.438$ $-1.504$ $-0.333$
Squirrel GCN $+0.006$ $-0.389$ $-8.183$ $-0.180$ $-0.066$ $+0.189$
GIN $+0.179$ $-0.204$ $+0.199$ $-0.052$ $-0.111$ $+0.085$
GAT $-0.393$ $-0.750$ $-1.899$ $-2.251$ $-1.165$ $-0.038$
TwitchDE GCN $-0.099$ $-0.113$ $-2.272$ $-0.426$ $-0.280$ $+0.165$
GIN $-0.079$ $-0.116$ $+0.140$ $-0.005$ $-0.389$ $-0.016$
GAT $-0.201$ $-1.206$ $-0.726$ $-0.938$ $-0.921$ $-0.129$
Mutag GCN $-0.087$ $-0.448$ $-0.069$ $-0.017$ $-0.026$ $+0.174$
GIN $-0.107$ $-0.351$ $-0.330$ $-0.640$ $-0.240$ $-0.342$
GAT $+0.071$ $+0.034$ $+0.099$ $+0.151$ $+0.091$ $+0.281$
Proteins GCN $+0.339$ $\mathbf{+0.450}$ $+0.260$ $+0.307$ $\mathbf{+0.397}$ $\mathbf{+0.477}$
GIN $-0.331$ $-0.541$ $-0.352$ $-0.491$ $-0.657$ $-0.217$
GAT $-0.076$ $-0.042$ $+0.040$ $+0.031$ $-0.136$ $-0.021$
Enzymes GCN $-0.072$ $-0.272$ $-0.055$ $-0.359$ $-0.758$ $-0.543$
GIN $-0.673$ $-0.070$ $-0.083$ $-0.548$ $-0.290$ $-0.048$
GAT $+0.031$ $-0.508$ $+0.055$ $-0.380$ $-0.960$ $-0.353$
Reddit GCN $-1.712$ $-1.333$ $-3.075$ $-2.360$ $-1.065$ $-1.527$
GIN $-0.531$ $+0.095$ $+0.163$ $+0.240$ $\mathbf{+0.503}$ $+0.165$
GAT $+0.209$ $\mathbf{+0.638}$ $+0.154$ $-0.054$ $\mathbf{+0.632}$ $\mathbf{+0.469}$
IMDb GCN $+0.246$ $\mathbf{+0.513}$ $\mathbf{+0.642}$ $+0.281$ $+0.200$ $+0.240$
GIN $-0.364$ $-0.024$ $-0.327$ $-0.932$ $-0.204$ $-0.090$
GAT $-0.030$ $\mathbf{+0.917}$ $+0.245$ $+0.344$ $\mathbf{+0.885}$ $\mathbf{+0.650}$
Collab GCN $-0.195$ $-0.859$ $-6.962$ $+0.140$ $-0.348$ $-0.065$
GIN $-0.220$ $\mathbf{+0.396}$ $-0.916$ $-1.139$ $-0.071$ $\mathbf{+0.396}$
GAT $+0.000$ $+0.000$ $+0.000$ $+0.000$ $+0.000$ $+0.000$
The reliance on p-values as a measure of statistical significance has been widely criticized due to its limitations in conveying the magnitude of an effect. Although a low p-value indicates that an observed difference is unlikely to have occurred under the null hypothesis, it does not provide information about the practical significance of the result. A statistically significant effect may be too small to be meaningful in real-world applications, while a non-significant result does not necessarily imply the absence of a meaningful effect, particularly when sample sizes are small. These concerns have led to an increased emphasis on effect size measures, which quantify the magnitude of differences independently of sample size.
One widely used measure of effect size is Cohen’s $d$ statistic [14], which standardizes the difference between two group means by dividing by the pooled standard deviation:
$d$ $=\frac{\bar{x}_{1}-\bar{x}_{2}}{s_{p}}$ (G.1)
$s_{p}$ $=\sqrt{\frac{\left(n_{1}-1\right)s_{1}^{2}+\left(n_{2}-1\right)s_{2}^{2}}{n_{1}+n_{2}-2}}$ (G.2)
where $\bar{x}_{1}$ and $\bar{x}_{2}$ are sample means, $s_{1}^{2}$ and $s_{2}^{2}$ are unbiased sample variance, and $n_{1}$ and $n_{2}$ are the sizes of the samples. However, Cohen’s $d$ assumes that the sample standard deviation is an unbiased estimator of the population standard deviation. In small samples, this assumption does not hold, as the sample standard deviation tends to underestimate the true population variability. To address this bias, Hedges’ $g$ statistic [39] introduces a correction factor that adjusts Cohen’s $d$ statistic for small sample sizes:
$g\approx\left(1-\frac{3}{4\left(n_{1}+n_{2}\right)-9}\right)d$ (G.3)
Cohen [14] suggested that an effect size of $0.2$ be considered small, $0.5$ be considered medium, and $0.8$ be considered large. In Table 7, we present Hedges’ $g$ statistic for the statistical tests in Section 5.2. We can clearly see that for homophilic datasets, there is a strong positive effect of using the dropping methods, but for heterophilic datasets and graph classification datasets, there is at most a small positive effect of using the dropping methods; rather, in most cases there is a negative effect.
G.3 Performance of GIN with DropSens
Table 8: Performance of GIN with different rewiring methods in graph-classification tasks, following [44, 8]. First, second, and third best results are colored.
Rewiring Mutag Proteins Enzymes Reddit IMDb Collab
None $77.70$ $70.80$ $33.80$ $86.79$ $70.18$ 72.99
Last FA 83.45 72.30 47.40 90.22 70.91 75.06
Every FA $72.55$ $70.38$ $28.38$ $50.36$ $49.16$ $32.89$
DIGL 79.70 $70.76$ 35.72 $76.04$ $64.39$ $54.50$
SDRF 78.40 $69.81$ 35.82 $86.44$ $69.72$ $72.96$
FoSR $78.00$ 75.11 $29.20$ 87.35 71.21 73.28
GTR $77.60$ 73.13 $30.57$ 86.98 71.28 $72.93$
DropSens $70.60$ $68.00$ $29.56$ $76.44$ $62.48$ $65.77$
In Section 5.3, we showed that when modelling long-range graph-classification tasks using GCNs, DropSens outperforms state-of-the-art graph-rewiring techniques designed for alleviating over-squashing. However, it does not perform as well with GIN, as can be seen in Table 8 – unsurprising, since DropSens was specifically designed to work with GCN’s message-passing scheme. This highlights the main limitation of DropSens, necessitating architecture-specific alteration to the edge-dropping strategy, which is not practical in general.
G.4 Best-performing Dropping Probabilities
Table 9: Best performing dropout configuration – $q_{\max}$ and $c$ for DropSens, and $q$ for other dropping methods.
GNN Dataset DropEdge DropNode DropAgg DropGNN Dropout DropMessage DropSens
Cora GCN $0.7$ $0.4$ $0.1$ $0.7$ $0.7$ $0.1$ $0.5$, $0.95$
GIN $0.1$ $0.2$ $0.1$ $0.1$ $0.3$ $0.5$ $0.2$, $0.90$
GAT $0.8$ $0.3$ $0.9$ $0.4$ $0.7$ $0.6$ $0.8$, $0.50$
CiteSeer GCN $0.9$ $0.1$ $0.9$ $0.8$ $0.1$ $0.3$ $0.2$, $0.95$
GIN $0.1$ $0.4$ $0.1$ $0.1$ $0.1$ $0.2$ $0.3$, $0.95$
GAT $0.9$ $0.2$ $0.9$ $0.8$ $0.1$ $0.1$ $0.8$, $0.50$
PubMed GCN $0.3$ $0.4$ $0.1$ $0.2$ $0.5$ $0.7$ $0.5$, $0.95$
GIN $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.6$ $0.2$, $0.90$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.5$ $0.8$ $0.3$, $0.95$
Chameleon GCN $0.4$ $0.1$ $0.1$ $0.1$ $0.4$ $0.5$ $0.5$, $0.80$
GIN $0.1$ $0.1$ $0.2$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.95$
Squirrel GCN $0.2$ $0.6$ $0.9$ $0.3$ $0.5$ $0.6$ $0.5$, $0.90$
GIN $0.2$ $0.5$ $0.2$ $0.5$ $0.3$ $0.7$ $0.2$, $0.95$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
TwitchDE GCN $0.3$ $0.2$ $0.9$ $0.1$ $0.1$ $0.6$ $0.3$, $0.95$
GIN $0.2$ $0.3$ $0.1$ $0.2$ $0.1$ $0.5$ $0.2$, $0.90$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.90$
Actor GCN $0.9$ $0.1$ $0.9$ $0.5$ $0.2$ $0.1$ $0.2$, $0.95$
GIN $0.5$ $0.1$ $0.2$ $0.3$ $0.6$ $0.2$ $0.2$, $0.90$
GAT $0.9$ $0.2$ $0.8$ $0.7$ $0.1$ $0.1$ $0.8$, $0.50$
Mutag GCN $0.5$ $0.8$ $0.7$ $0.2$ $0.2$ $0.3$ $0.5$, $0.80$
GIN $0.1$ $0.3$ $0.1$ $0.6$ $0.2$ $0.5$ $0.5$, $0.95$
GAT $0.4$ $0.1$ $0.2$ $0.1$ $0.4$ $0.4$ $0.5$, $0.95$
Proteins GCN $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.3$ $0.8$, $0.90$
GIN $0.1$ $0.3$ $0.1$ $0.1$ $0.4$ $0.1$ $0.5$, $0.95$
GAT $0.4$ $0.1$ $0.2$ $0.1$ $0.1$ $0.1$ $0.8$, $0.95$
Enzymes GCN $0.5$ $0.1$ $0.1$ $0.6$ $0.1$ $0.4$ $0.8$, $0.80$
GIN $0.3$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.8$, $0.95$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.5$, $0.90$
Reddit GCN $0.1$ $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.95$
GIN $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
GAT $0.9$ $0.7$ $0.6$ $0.7$ $0.6$ $0.9$ $0.8$, $0.95$
IMDb GCN $0.6$ $0.7$ $0.1$ $0.6$ $0.1$ $0.9$ $0.8$, $0.50$
GIN $0.1$ $0.3$ $0.1$ $0.1$ $0.2$ $0.2$ $0.2$, $0.95$
GAT $0.7$ $0.7$ $0.1$ $0.6$ $0.9$ $0.6$ $0.2$, $0.95$
Collab GCN $0.1$ $0.1$ $0.4$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
GIN $0.1$ $0.2$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.90$
GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.90$
For the real-world datasets in Section 5.2, we report the best performing dropping probability for different dataset--model--dropout combinations in Table 9.